the owners of candy store want to sell, for $6 per pound, a mixture of chocolate covered raisins, which usually sell for $3 per pound, and choclate covered macadamia nuts, which usually sells for $8 per pound, they have a 70 pound barrel of raisins . How much pounds of nuts should they mix with the barrel of raisin so they hit their target value of $6 per pound for the mixture








Answer by stanbon(75887)

the owners of candy store want to sell, for $6 per pound, a mixture of chocolate covered raisins, which usually sell for $3 per pound, and choclate covered macadamia nuts, which usually sells for $8 per pound, they have a 70 pound barrel of raisins . How much pounds of nuts should they mix with the barrel of raisin so they hit their target value of $6 per pound for the mixture
--------------
raisins = 70 lbs
nuts = n lbs
----
Equations:
Value: 3*70 + 8n = 6(70+n)
---
3*70 + 8n = 6*70 + 6n
2n = 3*70
n = 3*35 = 105 lbs (amout of nuts needed)
=============================
Cheers,
Stan H.
=============